# Revela: Dense Retriever Learning via Language Modeling

### v2

@@ Abstract
To answer this question, we introduce Revela, a unified and scalable training framework for self-supervised retriever learning via language modeling. Revela models semantic dependencies among documents by conditioning next token prediction on local and cross-document context through an in-batch attention mechanism. This attention is weighted by retriever-computed similarity scores, enabling the retriever to be optimized as part of language modeling. We evaluate Revela on domain-specific (CoIR), reasoning-intensive (BRIGHT), and general-domain (BEIR) benchmarks across various retriever backbones. Without annotated or synthetic query-document pairs, Revela surpasses larger supervised models and proprietary APIs on CoIR and matches them on BRIGHT. It achieves BEIR’s unsupervised SoTA with ~ 1000x less training data and 10x less compute. Performance increases with batch size and model size, highlighting Revela’s scalability and its promise for self‑supervised retriever learning.
@@ 1 Introduction
We comprehensively demonstrate the effectiveness of Revela across three benchmarks in § 4: CoIR (Li et al., 2024), a benchmark tailored for code retrieval, BRIGHT (Hongjin et al., 2025), a reasoning-intensive benchmark across diverse domains, and BEIR (Thakur et al., 2021), a heterogeneous benchmark covering general domains. To do this, we train Revela on Wikipedia for general retrieval and on code-related corpora (Wang et al., 2025) for code retrieval, using pretrained transformers ranging from 135M to 8B parameters as retriever backbones paired with an LM. On CoIR, Revela outperforms a strong, 7B-parameter supervised retriever (E5-Mistral-7b-Instruct) by 2.8 points and surpasses the weakly-supervised baseline E5-PT (Wang et al., 2022) by 9.7 points at a similar scale. This is particularly noteworthy as both baselines was pre-trained on massive query-document pairs that encompass Revela’s training data. Furthermore, Revela matches proprietary models on BRIGHT and achieves parity with weakly-supervised E5 model on BEIR while using approximately 1,000x less training data and 10x less compute. These results establish Revela as a highly effective and efficient self-supervised solution. In § 5, we also show that scaling Revela via on larger retriever backbones, larger LMs, and larger batch size can yield greater gain over baselines. Compared to traditional contrastive learning, Revela exhibits stronger cross-domain generalization. Moreover, the mixed-data training allows the model to scale across multiple domains while guaranteeing high domain-specific performance. Collectively, the evidence highlights Revela’s robust scaling behavior and strong generalization across models, data, and domains. To this end, we summarize our contributions as follows:
@@
  • Without query-document pairs, Revela surpasses E5-Mistral-7b-Instruct on CoIR by 2.8% and outperforms unsupervised baselines by 9.7% at the comparable scale, while also matching the performance of proprietary APIs on BRIGHT, a challenging, reasoning-intensive benchmark.
@@ 3.1 Training Objectives
$P_{R}(x_{l}^{i})=P_{\Phi,\Theta}\!\bigl(x_{l}^{i}\mid x_{<l}^{i},\{D_{j}\}_{j\neq i}\bigr)$ (2)
@@ 3.2 In-batch Attention
combines standard self-attention with cross-document attention. The embeddings of $k$-th token in $D_{i}$ is obtained by (1) the prefix of $D_{i}$, and (2) the other documents $\{D_{j}\}_{j\neq i}$ based on their similarity to $D_{i}$. This encourages $D_{i}$ to selectively attend to more relevant documents based on learned retrieval signals. For the contribution from the prefix of $D_{i}$, the self-attention output $\mathrm{s}_{i}^{l}$ in the in-batch attention uses the same projection matrices as above
@@
Cross-document attention enables $D_{i}$ to attend to other documents $D_{j}$ using cached keys $K_{j}^{e}$ and values $V_{j}^{e}$. With a full attention mask, the output $\mathrm{b}^{l}_{ij}$ is computed as
@@
$\mathrm{h}^{l}_{i}=\mathrm{s}^{l}_{i}+\mathrm{b}^{l}_{i}$ (7)
@@ 3.3 Similarity Computation
The resulting weights $\mathrm{Sim}(D_{i},D_{j})\in\mathbb{R}^{B\times B}$ capture cross-document similarities and are used to condition language modeling on relevant in-batch documents as presented above.
@@ 4.1 Experimental Setups
To comprehensively evaluate our proposed framework, we benchmark Revela on three diverse datasets: CoIR (Li et al., 2024), a comprehensive benchmark designed for code-specific retrieval tasks, BRIGHT (Hongjin et al., 2025), a retrieval benchmark requiring intensive reasoning to retrieve relevant documents spanning diverse domains, and BEIR (Thakur et al., 2021), a heterogeneous benchmark covering multiple domains for general information retrieval. A more detailed introduction is listed in § B.1.
@@
  • BRIGHT & BEIR: Similarly, we segment the passages in the Wikipedia corpus [^1] https://huggingface.co/datasets/Tevatron/wikipedia-nq-corpus. Given a set of passages $\{d_{1},d_{2},\ldots,d_{n}\}$, where each passage $d_{i}$ is divided into chunks $(d_{i1},d_{i2},\ldots,d_{im_{i}})$, the chunks are interleaved in the order $(d_{11},d_{21},\ldots,d_{n1},d_{12},d_{22},\ldots)$ and then grouped sequentially into batches of size 16. In total, we sample 320,000 batches from 339,409 documents for training. Notably, a single batch may contain chunks from different documents, highlighting the flexibility of batch construction in Revela.
@@
Revela jointly trains a retriever and an LM using the NTP objective. We adopt LLaMA-3.2-1B (Grattafiori et al., 2024) as the LM. To ensure a fair comparison across diverse unsupervised baselines, we adopt a range of LMs with parameter sizes from 0.1B to 3B as the backbone models for the retrievers: SmolLM2-135M (Allal et al., 2025), Qwen2.5-0.5B (Yang et al., 2024), LLaMA-3.2-1B and LLaMA-3.2-3B. [^2] For computational efficiency, we set smaller batch size for the 3B and 8B model, e.g., 8. For BRIGHT, we specifically include LLaMA-3.1-8B as the retriever’s backbone for the comparison with large LM-based and proprietary retrievers. We follow the approach used in RepLLaMA (Ma et al., 2024), appending the <eos> token to each sentence and use its corresponding embedding as the sentence representation. Additionally, we prepend the prefixes ”Query: ” and ”Passage: ” to queries and passages, respectively. For more details about model checkpoints, please refer to § B.4.
@@
We include representative self-supervised baselines: E5-PTlarge (E5-PT; Wang et al. 2022) is trained with a contrastive objective, leveraging weak supervision signals from a curated large-scale dataset of text pairs (1.3B raw pairs, filtered to 270M) spanning multiple domains (e.g., code) and covering Revela’s training corpus. REPLUG (Shi et al., 2024) distills supervision from a frozen LM into a retriever by using LM perplexity to model within-batch chunk–chunk similarity, conditioning one chunk on the other. We adopt REPLUG [^3] For replication details, see § B.3. as our main baseline because (1) it matches Revela’s retriever as decoder-only LM design, unlike encoder–decoder systems such as Atlas (Izacard et al., 2023), keeping the focus on joint retriever–LM training; (2) this architectural match makes comparisons generalizable across scales; and (3) REPLUG outperforms most prior methods, making it representative of this line of work (Shi et al., 2024).
For CoIR, we also include several supervised retrievers, as well as API-based models such as OpenAI-Ada-002 (Ada-2; Neelakantan et al. 2022) and Voyage-Code-002 (Voyage-2; Voyage AI 2024), following the original setup (Li et al., 2024). Supervised retrievers include UniXcoder (UniX; Guo et al. 2022), which is finetuned on code-related datasets; BGE-M3 (BGE; Chen et al. 2024), a supervised model pretrained and finetuned on text pairs (including code); and E5-Mistral-7B-Instruct (E5-Mistral; Wang et al. 2024).
For BRIGHT, we use several strong baselines: API-based models from text-embedding-3-large (OpenAI), cohere-embed-english-v3.0 (Cohere; Cohere 2023), and voyage-large-2-instruct (VoyageAI), as well as E5-Mistral.
@@
For more details about the Huggingface checkpoints, please refer to § B.4.
@@
For both retrievers and LMs, we apply LoRA with a rank of 256. Training uses the temperature $\tau$ $1\mathrm{e}{-4}$, a learning rate of $1\mathrm{e}{-4}$, and 100 warmup steps, following the WarmupDecayLR schedule. We train on 4× A100 80GB GPUs with a gradient accumulation step size of 8. Passages are truncated to 160 tokens, and bf16 mixed precision is enabled. We finetune the models for one epoch, namely, there are 10,000 steps on Wikipedia (~ 44 hours) and around 11,000 steps on code-related texts (~ 48 hours). During inference, the max token length of queries and documents is 2048. For more details of the experimental setups, please refer to § B.5.
@@ 4.2 Experimental Results
Tab. 1 reports the performance of Revela and baseline methods on CoIR. As an unsupervised model trained without query-document pairs, Revela3B surpasses E5-Mistral-7B-Instruct, a much larger supervised model pre-trained and fine-tuned on massive, well-curated text-doc pairs, as well as two proprietary APIs averaged on 10 tasks. Revela also follows scaling laws: its performance consistently improves with model size, while maintaining superiority at every scale over baselines. At 0.1B parameters, Revela outperforms the code-specific supervised model UniXCoder by 11.1 points on NDCG@10. At the 0.5B scale, our model outperforms E5-PT by nearly 10 points, despite E5-PT being pre-trained on 270 million filtered query-document pairs covering Revela’s corpus. Revela0.5B even surpasses the supervised model BGE-M3, despite the latter being pre-trained on extensive text-code pairs. [^4] Please refer to Table 8 in the original paper (Chen et al., 2024). Moreover, at each scale, Revela also surpasses REPLUG, underscoring the effectiveness of the co-training paradigm between retrievers and LMs.
@@
Figure 3: Performance on BRIGHT (Left) and BEIR (Right) (NDCG@10,%). Results for Revelaare shown in opaque bars, while all other models are represented by transparent bars. On BRIGHT, a reasoning-intensive task, Revela performs on par with large supervised models and properties APIs. On BEIR, Revela achieves similar performance with E5-PT with much less data and compute. Please refer to Tab. 6 and Tab. 7 in § B.6 for the per-task results.
@@
Fig. 3 (Left) shows the average NDCG@10 across the 12 BRIGHT subtasks. Trained unsupervised on only 340K raw Wikipedia documents, Revela8B performs on par with the supervised E5-Mistral-7B-Instruct and other proprietary models, and even surpasses cohere-embed-english-v3.0. The scaling trends observed in code retrieval persist here as well: even Revela0.5B outperforms REPLUG3B with much fewer parameters. At a comparable parameter scale, Revela0.5B exceeds E5-PT by 1.1 points (10% relative), a noteworthy result given the potential advantage from E5-PT’s training corpus overlapping with BRIGHT. These findings highlight Revela’s promise for tackling more complex retrieval scenarios.
@@
Fig. 3 (Right) reports the average NDCG@10 across the 13 BEIR tasks, where Revela proves to be remarkably efficient and robust. At the 0.1B scale, Revela outperforms Contriever and LaPraDor by more than 3 absolute points. Moreover, Revela’s consistently surpass REPLUG by a significant margin (3B: 8.9%; 1B: 7.0%), mirroring trends observed on CoIR and BRIGHT. Remarkably, despite requiring roughly 1000× less training data and over 10× fewer compute resources, Revela3B achieves performance on par with E5-PT, underscoring its efficiency and effectiveness. Notably, even when trained solely on a code-related corpus, Revela3B performs comparably on the general-domain BEIR benchmark to both its Wikipedia-trained counterpart and E5-PT, underscoring its powerful generalization capabilities.
@@ 5 Analysis
We conduct several targeted analyses to further investigate Revela. First, to isolate its algorithmic contribution, we compare it with Contriever using an identical LM backbone. This comparison reveals that Revelaachieves superior performance and exhibits stronger domain-specific robustness to training data. Second, we demonstrate that, consistent with traditional contrastive learning, Revelabenefits from larger batch sizes. Finally, we examine the LM’s impact on retriever performance within the co-training framework. Additional studies on mixed-domain training, out-of-domain generalization, the LM’s post-Revela capabilities, and computational efficiency are presented in Appendix App. C.
@@
As shown in Fig. 5, CoIR exhibits a clear positive trend: the largest LM delivers best retrieval performance. By contrast, on BEIR, the general-domain benchmark, larger LMs do not provide a consistent advantage. We therefore expect that seamlessly incorporating larger LMs will further improve Revela’s adaptability to specialized domains. For per-dataset results, see § C.3.

### v3

@@ Abstract
To answer this question, we introduce Revela, a unified and scalable training framework for self-supervised retriever learning via language modeling. Revela models semantic dependencies among documents by conditioning next token prediction on local and cross-document context through an in-batch attention mechanism. This attention is weighted by retriever-computed similarity scores, enabling the retriever to be optimized as part of language modeling. We evaluate Revela on domain-specific (CoIR), reasoning-intensive (BRIGHT), and general-domain (BEIR) benchmarks across various retriever backbones. Without annotated or synthetic query-document pairs, Revela surpasses larger supervised models and proprietary APIs on both CoIR and BRIGHT. It achieves BEIR’s unsupervised SoTA with ~ 1000x less training data and 10x less compute. Performance increases with batch size and model size, highlighting Revela’s scalability and its promise for self‑supervised retriever learning.
@@ 1 Introduction
We comprehensively demonstrate the effectiveness of Revela across three benchmarks in § 4: CoIR (Li et al., 2025), a benchmark tailored for code retrieval, BRIGHT (Hongjin et al., 2025), a reasoning-intensive benchmark across diverse domains, and BEIR (Thakur et al., 2021), a heterogeneous benchmark covering general domains. To do this, we train Revela on Wikipedia for general retrieval and on code-related corpora (Wang et al., 2025) for code retrieval, using pretrained transformers ranging from 135M to 3B parameters as retriever backbones paired with an LM. On CoIR, Revela outperforms a strong, 7B-parameter supervised retriever (E5-Mistral-7b-Instruct) by 2.8 points (nDCG@10) and surpasses the weakly-supervised baseline E5-PT (Wang et al., 2022) by 9.7 points at a similar scale. This is particularly noteworthy as both baselines was pre-trained on massive query-document pairs that encompass Revela’s training data. Furthermore, Revela outperforms proprietary models on BRIGHT and achieves parity with weakly-supervised E5 model on BEIR while using approximately 1,000x less training data and 10x less compute. These results establish Revela as a highly effective and efficient self-supervised solution. In § 5, we also show that scaling Revela via on larger retriever backbones, larger LMs, and larger batchs can yield greater gain over baselines. Compared to conventional contrastive learning, Revela exhibits stronger cross-domain generalization. Moreover, the mixed-data training allows the model to scale across multiple domains while guaranteeing high domain-specific performance. Collectively, the evidence highlights Revela’s robust scaling behavior and strong generalization across models, data, and domains. To this end, we summarize our contributions as follows:
@@
  • Without query-document pairs, Revela surpasses E5-Mistral-7b-Instruct on CoIR by 2.8% (nDCG@10) and outperforms unsupervised baselines by 9.7% at the comparable scale, while also outperforming proprietary APIs on BRIGHT, a challenging, reasoning-intensive benchmark.
@@ 3.1 Training Objectives
$P_{R}(x_{l}^{i})=P_{\Phi,\Theta}\!\bigl(x_{l}^{i}\mid x_{<l}^{i},\{D_{j}\}_{j\neq i}\bigr).$ (2)
@@ 3.2 In-batch Attention
combines standard self-attention with cross-document attention. The embeddings of $k$-th token in $D_{i}$ is obtained by (1) the prefix of $D_{i}$, and (2) the other documents $\{D_{j}\}_{j\neq i}$ based on their similarity to $D_{i}$. This encourages $D_{i}$ to selectively attend to more relevant documents based on learned retrieval signals.
For the contribution from the prefix of $D_{i}$, the self-attention output [^1] Note that the self-attention output $\mathrm{s}_{i}^{l}$ is distinct from the computation of $\mathrm{e}_{i}^{l}$. As shown in Fig. 2, $\mathrm{e}_{i}^{l}$ and $\mathrm{s}_{i}^{l}$ correspond to the upper-left and bottom-right corners of the attention map, respectively. Functionally, $\mathrm{e}_{i}^{l}$ captures only sequence-level information, whereas $\mathrm{s}_{i}^{l}$ contains in-batch information derived from $\mathrm{h}_{i}^{l-1}$. $\mathrm{s}_{i}^{l}$ in the in-batch attention uses the same projection matrices as above
@@
Cross-document attention enables $D_{i}$ to attend to other documents $D_{j}$ using cached keys $K_{j}^{e}$ and values $V_{j}^{e}$. With a full attention mask, the bottom-left corner in Fig. 2, the output $\mathrm{b}^{l}_{ij}$ is computed as
@@
$\mathrm{h}^{l}_{i}=\mathrm{s}^{l}_{i}+\mathrm{b}^{l}_{i}.$ (7)
@@ 3.3 Similarity Computation
The resulting pairwise weights $\mathrm{Sim}(D_{i},D_{j})]_{i,j=1}^{B}\in\mathbb{R}^{B\times B}$ capture cross-document similarities, allowing the model to condition on relevant in-batch documents.
@@ 4.1 Experimental Setups
To comprehensively evaluate our proposed framework, we benchmark Revela on three diverse datasets: CoIR (Li et al., 2025), a comprehensive benchmark designed for code-specific retrieval tasks, BRIGHT (Hongjin et al., 2025), a retrieval benchmark requiring intensive reasoning to retrieve relevant documents spanning diverse domains, and BEIR (Thakur et al., 2021), a heterogeneous benchmark covering multiple domains for general information retrieval. A more detailed introduction is listed in § B.1.
@@
  • BRIGHT & BEIR: Similarly, we segment the passages in the Wikipedia corpus. [^2] https://huggingface.co/datasets/Tevatron/wikipedia-nq-corpus Given a set of passages $\{d_{1},d_{2},\ldots,d_{n}\}$, where each passage $d_{i}$ is divided into chunks $(d_{i1},d_{i2},\ldots,d_{im_{i}})$, the chunks are interleaved in the order $(d_{11},d_{12},\ldots,d_{1m_{1}},d_{21},d_{22},\ldots)$ and then grouped sequentially into batches of size 16. In total, we sample 320,000 batches from 339,409 documents for training. Notably, a single batch may contain chunks from different documents, highlighting the flexibility of batch construction in Revela.
@@
Revela jointly trains a retriever and an LM using the NTP objective. We adopt LLaMA-3.2-1B (Grattafiori et al., 2024) as the LM. To ensure a fair comparison across diverse unsupervised baselines, we adopt a range of LMs with parameter sizes from 0.1B to 3B as the backbone models for the retrievers: SmolLM2-135M (Allal et al., 2025), Qwen2.5-0.5B (Yang et al., 2024), LLaMA-3.2-1B and LLaMA-3.2-3B. [^3] For computational efficiency, we set smaller batch size for the 3B model, e.g., 8. We follow the approach used in RepLLaMA (Ma et al., 2024), appending the <eos> token to each sentence and use its corresponding embedding as the sentence representation. Additionally, we prepend the prefixes ”Query: ” and ”Passage: ” to queries and passages, respectively. For more details about model checkpoints, please refer to § B.4.
@@
We include representative self-supervised baselines: E5-PTlarge (E5-PT; Wang et al. 2022) is trained with a contrastive objective, leveraging weak supervision signals from a curated large-scale dataset of text pairs (1.3B raw pairs, filtered to 270M) spanning multiple domains (e.g., code) and covering Revela’s training corpus. REPLUG (Shi et al., 2024) distills supervision from a frozen LM into a retriever by using LM perplexity to model within-batch chunk–chunk similarity, conditioning one chunk on the other. We adopt REPLUG [^4] For replication details, see § B.3. as our main baseline because (1) it matches Revela’s retriever as decoder-only LM design, unlike encoder–decoder systems such as Atlas (Izacard et al., 2023), keeping the focus on joint retriever–LM training; (2) this architectural match makes comparisons generalizable across scales; and (3) REPLUG outperforms most prior methods, making it representative of this line of work (Shi et al., 2024). Consistent with Revela, we use LLaMA-3.2-1B as the frozen reference LM in REPLUG.
For CoIR, we include several supervised retrievers, as well as API-based models such as OpenAI-Ada-002 (Ada-2) and Voyage-Code-002 (Voyage-2), following the original setup (Li et al., 2025). Supervised retrievers include UniXcoder (UniX; Guo et al. 2022), which is finetuned on code-related datasets; BGE-M3 (BGE; Chen et al. 2024), a supervised model pretrained and finetuned on text pairs (including code); and E5-Mistral-7B-Instruct (E5-Mistral; Wang et al. 2024).
For BRIGHT, we use several strong baselines: API-based models from text-embedding-3-large (OpenAI), cohere-embed-english-v3.0 (Cohere), and voyage-large-2-instruct (VoyageAI), as well as E5-Mistral.
@@
For more details about the model access and the Huggingface checkpoints, please refer to § B.4.
@@
For both retrievers and LMs, we apply LoRA with a rank of 256. Training uses the temperature $\tau$ $1\mathrm{e}{-4}$, a learning rate of $1\mathrm{e}{-4}$, and 100 warmup steps, following the WarmupDecayLR schedule. We train on 4 A100 80GB GPUs with a gradient accumulation step size of 8. Passages are truncated to 160 tokens, and bf16 mixed precision is enabled. We finetune the models for one epoch, namely, there are 10,000 steps on Wikipedia (~ 44 hours) and around 11,000 steps on code-related texts (~ 48 hours). During inference, the max token length of queries and documents is 2048. For more details of the experimental setups, please refer to § B.5.
@@ 4.2 Experimental Results
Tab. 1 reports the performance of Revela and baseline methods on CoIR. As an unsupervised model trained without query-document pairs, Revela3B surpasses E5-Mistral-7B-Instruct, a much larger supervised model pre-trained and fine-tuned on massive, well-curated query-doc pairs, as well as two proprietary APIs averaged on 10 tasks. Revela also follows scaling laws: its performance consistently improves with larger model sizes, while maintaining superiority at every scale over baselines. At 0.1B parameters, Revela outperforms the code-specific supervised model UniXCoder by 11.1 points on nDCG@10. At the 0.5B scale, our model outperforms E5-PT by nearly 10 points, despite E5-PT being pre-trained on 270 million filtered query-document pairs covering Revela’s corpus. Revela0.5B even surpasses the supervised model BGE-M3, despite the latter being pre-trained on extensive text-code pairs. [^5] Please refer to Table 8 in the original paper (Chen et al., 2024). Moreover, at each scale, Revela also surpasses REPLUG, underscoring the effectiveness of the retriever-LM co-training paradigm for retriever learning.
@@
Figure 3: Performance on BRIGHT (left) and BEIR (right) (nDCG@10, %). Results for Revela are shown in opaque bars, while all other models are represented by transparent bars. On BRIGHT, Revela3B surpasses E5-Mistral, a supervised retriever with more parameters, and properties APIs. On BEIR, Revela achieves similar performance with E5-PT with much less data and compute. Please refer to Tab. 7 and Tab. 8 in § B.6 for the per-task results.
@@
Fig. 3 (left) shows the average nDCG@10 across the 12 BRIGHT subtasks. Despite being trained unsupervised on only 340K raw Wikipedia documents, Revela3B outperforms the supervised E5-Mistral-7B-Instruct as well as proprietary embedding models. The scaling trends observed in code retrieval persist here as well: even Revela0.5B outperforms REPLUG3B with much fewer parameters. At comparable scale, Revela0.5B exceeds E5-PT by 3.1 points (23.7% relative), a noteworthy result given the potential advantage from E5-PT’s training corpus overlapping with BRIGHT. These findings highlight Revela’s promise for tackling more complex retrieval scenarios.
@@
Fig. 3 (right) reports the average nDCG@10 across the 13 BEIR tasks, where Revela demonstrates remarkable efficiency and robustness. At the 0.1B scale, Revela outperforms Contriever and LaPraDor by over 3 absolute points. Moreover, Revela’s consistently surpass REPLUG by a significant margin (3B: 8.9%; 1B: 7.0%), mirroring trends observed on CoIR and BRIGHT. Remarkably, despite using approximately 1000× less training data and 10× fewer compute resources, Revela3B matches E5-PT’s performance, underscoring its efficiency. Most notably, when trained solely on a code-related corpus, Revela3B performs comparably on the general-domain BEIR benchmark to both its Wikipedia-trained counterpart and E5-PT, demonstrating strong cross-domain generalization.
@@ 5 Analysis
We conduct several targeted analyses to further investigate Revela. First, to isolate its algorithmic contribution, we compare it with Contriever using an identical LM backbone. This comparison reveals that Revela achieves superior performance and exhibits stronger domain-specific robustness to training data. Second, we demonstrate that, consistent with traditional contrastive learning, Revela  benefits from larger batch sizes. Finally, we examine the LM’s impact on retriever performance within the co-training framework. Additional studies on mixed-domain training, out-of-domain generalization, the LM’s post-Revela capabilities, and computational efficiency are presented in App. C.
@@
As shown in Fig. 5, CoIR exhibits a clear positive trend: the largest LM delivers best retrieval performance. In contrast, on the general-domain BEIR benchmark, larger LMs do not provide a consistent advantage. These findings suggest that incorporating larger LMs will likely enhance Revela’s performance on specialized domains while maintaining competitive results on general-domain tasks. For per-dataset results, see § C.3.